OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of ?

OpenStudy (zzr0ck3r):

y = mx+b but y = 3 and x = 6 3=6m+b b = 3-6m = 3(1-2m) so y = mx + 3(1-2m)